# On Spheres with $k$ Points Inside

### v2

@@ 2.3 The Finite Case
Let $A\subseteq{{\mathbb{R}}}^{d}$ be finite and generic, $k\geq 0$ an integer, and $x\in{{\mathbb{R}}}^{d}$ generic with respect to $A$. Then $x$ is covered by at most $\binom{d+k}{d}$ $k$-hefty simplices of $A$, with equality iff $x\in{{H}_{k}{({A})}}$.
@@
To prove the upper bound, we add points outside all circumspheres of $d+1$ points in $A$ to construct a thin Delaunay set $A^{\prime}\subseteq{{\mathbb{R}}}^{d}$. This is possible because the union of balls bounded by such spheres is bounded. Hence, $A\subseteq A^{\prime}$, and any $k$-hefty simplex of $A$ is also a $k$-hefty simplex of $A^{\prime}$. Assuming $x$ is generic with respect to $A^{\prime}$, Theorem 2.3 implies that exactly $\binom{d+k}{d}$ $k$-hefty simplices of $A^{\prime}$ cover $x$. If all of them are also $k$-hefty simplices of $A$, then $x$ is covered by exactly $\binom{d+k}{d}$ $k$-hefty simplices of $A$, but if not, then $x$ is covered by fewer than $\binom{d+k}{d}$ $k$-hefty simplices of $A$.
@@
$\Longrightarrow$”. Assuming $x\not\in{{H}_{k}{({A})}}$, we modify the construction of $A^{\prime}$ to show that there is a set $A^{\prime\prime}\supseteq A$ such that all $k$-hefty simplices of $A$ are $k$-hefty simplices of $A^{\prime\prime}$ but at least one $k$-hefty simplex of $A^{\prime\prime}$ that covers $x$ is not a $k$-hefty simplex of $A$. Since $x\not\in{{H}_{k}{({A})}}$, there are $d$ points in $A$ such that the open half-space bounded by the hyperplane passing through these points that contains $x$ contains at most $k$ points of $A$. Let $\Delta$ be the $(d-1)$-simplex spanned by the $d$ points, write $\Delta^{+}$ for the open half-space, and let $B\subseteq A$ be the points in $\Delta^{+}$. By assumption, #${{\#}{B}}\leq k$. We get $A^{\prime\prime}$ by adding $1+k-{{\#}{B}}$ points to $A$ as follows. First, we add $k-{{\#}{B}}$ points in $\Delta^{+}$ but outside all circumspheres of $d+1$ points in $A$. After that, we add a point $y\in{{\mathbb{R}}}^{d}$ so that the $d$-simplex that is the pyramid with apex $y$ and base $\Delta$ covers $x$, and all other $k$ points in $\Delta^{+}$ are inside the circumsphere of this $d$-simplex. We also require that $y$ is outside all circumspheres of $d+1$ points in $A$. It is clear that such a point $y$ exists on a sufficiently large sphere that passes through all vertices of $\Delta$. As proved above, there are at most $\binom{d+k}{d}$ $k$-hefty simplices of $A^{\prime\prime}$ that cover $x$. The pyramid with apex $y$ and base $\Delta$ is such a simplex, but it is not a $k$-hefty simplex of $A$. Hence, the number of $k$-hefty simplices of $A$ that cover $x$ is strictly less than $\binom{d+k}{d}$, as claimed.
2.4 Local Covering
By Theorem 2.3, the $k$-hefty simplices of a thin Delaunay set cover ${{\mathbb{R}}}^{d}$ without gap an integer number of times. However, beyond one dimension, it is generally not possible to split such a cover into sub-covers (separate layers) that enjoy the same property. Such splits are however possible locally, even in neighborhoods of the vertices of the simplices.
@@ 2.4 Local Covering
Let $A\subseteq{{\mathbb{R}}}^{d}$ be a generic thin Delone set and $k$ a non-negative integer. Then the $k$-hefty simplices of $A$ that share a vertex $a\in A$ cover any sufficiently small neihgbourhood of $a$ exactly $\binom{d+k-1}{d-1}$ times.
Proof 2.8.
The case $k=0$ is that of the Delaunay triangulation of $A$. Its top-dimensional simplices cover ${{\mathbb{R}}}^{d}$ exactly once, and thus also every neighborhood of $a\in A$ exactly once. We therefore assume $k>0$. Let $A^{\prime}=A\setminus\{a\}$. Observe that $A^{\prime}$ is also a generic thin Delone set and $a$ is generic with respect to $A^{\prime}$. We can therefore apply Theorem 2.3 to the $(k-1)$-hefty simplices of $A^{\prime}$ that contain $a\in{{\mathbb{R}}}^{d}$. There are only finitely many such simplices and each of them contains a small neighborhood of $a$.
Let $y$ be a point in a sufficiently small neighborhood of $a$ that is inside the intersection of all $(k-1)$-hefty simplices of $A^{\prime}$ that contain $a$, and assume that $y$ is generic with respect to $A$. By Theorem 2.3, there are exactly $\binom{d+k}{d}$ $k$-hefty simplices of $A$ that cover $y$. They split into simplices incident and not incident to $a$. The latter are the $(k-1)$-hefty simplices of $A^{\prime}$ that contain $a$. Indeed, since $y$ is sufficiently close to $a$, a $k$-hefty simplex of $A$ with vertices in $A^{\prime}$ contains $a$ in its interior iff it contains $y$ in its interior. The circumsphere of each such simplex encloses exactly $k-1$ points of $A^{\prime}$ as we removed $a$ from $A$ to get $A^{\prime}$. This implies that the number of $k$-hefty simplices of $A$ that contain $y$ and are not incident to $a$ is $\binom{d+k-1}{d}$. Thus, the number of $k$-hefty simplices of $A$ that contain $y$ and are incident to $a$ is
$\binom{d+k}{d}-\binom{d+k-1}{d}=\binom{d+k-1}{d-1},$ (1)
as claimed.
Similar to Theorem 2.5, we establish an inequality for finite sets that follows from Theorem 2.7. We are satisfied with a less refined statement than formulated for the global case in Section 2.3.
Corollary 2.9.

### v3

@@ 2.3 Local Covering
Let $A\subseteq{{\mathbb{R}}}^{d}$ be a generic thin Delone set, $k$ a non-negative integer, and $a\in A$ a point in this set. Then the $k$-hefty simplices of $A$ incident to $a$ cover any sufficiently small neighborhood of $a$ exactly $\binom{d+k-1}{d-1}$ times.
@@
The case $k=0$ is that of the Delaunay triangulation of $A$. Its top-dimensional simplices cover ${{\mathbb{R}}}^{d}$ exactly once, and thus also every neighborhood of $a\in A$ exactly once. We therefore assume $k>0$. Let $A^{\prime}=A\setminus\{a\}$. Observe that $A^{\prime}$ is also a generic thin Delone set and $a$ is generic with respect to $A^{\prime}$. We can therefore apply Theorem 2.3 to the (k1)(k-1)-hefty simplices of $A^{\prime}$ that contain $a\in{{\mathbb{R}}}^{d}$. By Theorem 2.3, there are $\binom{d+k-1}{d}$ such simplices and each of them contains a small neighborhood of $a$.
Let $y$ be a point in a sufficiently small neighborhood of $a$ that is inside the intersection of the $\binom{d+k-1}{d}$ (k1)(k-1)-hefty simplices of $A^{\prime}$ containing $a$, and assume that $y$ is generic with respect to $A$. There are exactly $\binom{d+k}{d}$ $k$-hefty simplices of $A$ that contain $y$. These $\binom{d+k}{d}$ simplices split into those incident and not incident to $a$. Since $y$ is sufficiently close to $a$, the latter are the (k1)(k-1)-hefty simplices of $A^{\prime}$ that contain $a$. The circumscribed sphere of each such simplex encloses exactly $k$ points of $A$ but only k1k-1 points of $A^{\prime}$, which implies that the number of $k$-hefty simplices of $A$ that contain $y$ and are not incident to $a$ is $\binom{d+k-1}{d}$. It follows that the number of $k$-hefty simplices of $A$ that contain $y$ and are incident to $a$ is $\binom{d+k}{d}-\binom{d+k-1}{d}=\binom{d+k-1}{d-1}$, as claimed.
2.4 Finitely Many Points
For a finite set, $A\subseteq{{\mathbb{R}}}^{d}$, the $k$-th covering number predicted by Theorem 2.3 applies sufficiently deep inside the set but acts only as an upper bound near the fringes. To formalize this claim, we introduce a parametrized generalization of the convex hull: for each integer $k\geq 0$, the $k$-hull of $A$, denoted ${H}_{k}{({A})}$}, is the common intersection of all closed half-spaces that miss at most $k$ of the points in $A$. Clearly, ${{H}_{0}{({A})}}={{\rm conv\,}{A}}$, and ${{H}_{k}{({A})}}\supseteq{{H}_{k+1}{({A})}}$ for every $k$. It is also easy to see that ${H}_{k}{({A})}$} is contained in the convex hull of $A$ after removing at most $k$ of its points, and indeed ${H}_{k}{({A})}$} is the intersection of all such convex hulls. By the Centerpoint Theorem of discrete geometry [9, Section 4.1], ${H}_{k}{({A})}$$\neq\emptyset$ if $k<\frac{n}{d+1}$, in which n=#An={{\#}{A}} is the number of points in $A$.
Theorem 2.7.
Let $A\subseteq{{\mathbb{R}}}^{d}$ be finite and generic, $k\geq 0$ an integer, and $a\in A$. Then every $x\in{{\mathbb{R}}}^{d}$ that is generic with respect to $A$ is covered by at most $\binom{d+k}{d}$ $k$-hefty simplices of $A$, with equality iff $x\in{{H}_{k}{({A})}}$, and if $x$ is sufficiently close to $a$, then it is covered by at most $\binom{d+k-1}{d-1}$ $k$-hefty simplices incident to $a$.
Proof 2.8.
To prove the two upper bounds, we add points outside all circumscribed spheres of $d+1$ points in $A$ to construct a thin Delone set $A^{\prime}\subseteq{{\mathbb{R}}}^{d}$. This is possible because the union of balls bounded by such spheres is bounded. Hence, $A\subseteq A^{\prime}$, and any $k$-hefty simplex of $A$ is also a $k$-hefty simplex of $A^{\prime}$. Assuming $x$ is generic with respect to $A^{\prime}$, Theorems 2.3 and 2.5 imply that exactly $\binom{d+k}{d}$ $k$-hefty simplices of $A^{\prime}$ cover $x$, and exactly $\binom{d+k-1}{d-1}$ are incident to $a$. Since some of the $k$-hefty simplices of $A^{\prime}$ may not be $k$-hefty simplices of $A$, these two numbers are only upper bounds. We continue by proving that $k$-th covering number is the correct number of $k$-hefty simplices of $A$ that contain $x$ iff $x\in{{H}_{k}{({A})}}$.
@@ 2.4 Finitely Many Points
$\Longrightarrow$”. Assuming $x\not\in{{H}_{k}{({A})}}$, we modify the construction of $A^{\prime}$ to show that there is a set $A^{\prime\prime}\supseteq A$ such that all $k$-hefty simplices of $A$ are $k$-hefty simplices of $A^{\prime\prime}$ but at least one $k$-hefty simplex of $A^{\prime\prime}$ that covers $x$ is not a $k$-hefty simplex of $A$. In contrast to $A^{\prime}$, the set $A^{\prime\prime}$ will be finite. Since $x\not\in{{H}_{k}{({A})}}$, there are $d$ points in $A$ such that the open half-space bounded by the hyperplane passing through these points that contains $x$ contains at most $k$ points of $A$. Let $\Delta$ be the (d1)(d-1)-simplex spanned by the $d$ points, write $\Delta^{+}$ for the open half-space that contains $x$, and let $B\subseteq A$ be the points in $\Delta^{+}$. By assumption, ${{\#}{B}}\leq k$. We get $A^{\prime\prime}$ by adding 1+k#B1+k-{{\#}{B}} points to $A$ as follows. First, we add k#Bk-{{\#}{B}} points in $\Delta^{+}$ but outside all circumscribed spheres of $d+1$ points in $A$. After that, we add a point $y\in{{\mathbb{R}}}^{d}$ so that the $d$-simplex that is the pyramid with apex $y$ and base $\Delta$ contains $x$, and all other $k$ points in $\Delta^{+}$ are inside the circumscribed sphere of this $d$-simplex. We also require that $y$ is outside all circumscribed spheres of $d+1$ points in $A$. It is clear that such a point $y$ exists on a sufficiently large sphere that passes through all vertices of $\Delta$. As proved above, there are at most $\binom{d+k}{d}$ $k$-hefty simplices of $A^{\prime\prime}$ that cover $x$. The pyramid with apex $y$ and base $\Delta$ is such a simplex, but it is not a $k$-hefty simplex of $A$. Hence, the number of $k$-hefty simplices of $A$ that cover $x$ is strictly less than $\binom{d+k}{d}$, as claimed.
[…]
2.5 Weighted Points
This section extends the results on covering with $k$-hefty simplices from points to weighted points in ${{\mathbb{R}}}^{d}$. This more general setting has a long history within mathematics and appears among other topics in the study of packings and coverings [19]; see [4] for a survey.
To introduce the formalism, we let ${\bf A}\subseteq{{\mathbb{R}}}^{d}\times{{\mathbb{R}}}$ be a locally finite subset of ${{\mathbb{R}}}^{d}$ in which each point is assigned a real weight. We use notation to distinguish a weighted point, ${\bf a}=(a,w_{a})$, from its location, $a\in{{\mathbb{R}}}^{d}$, and we write $w_{a}\in{{\mathbb{R}}}$ for its weight. The power or weighted squared distance of a point $x\in{{\mathbb{R}}}^{d}$ from $\bf a$} is ${\pi_{a}{({x})}}={\|{x}-{a}\|}^{2}-w_{a}$. A convenient geometric picture of $\bf a$} is the sphere with center $a\in{{\mathbb{R}}}^{d}$ and radius $r_{a}=\sqrt{w_{a}}$, which is either positive real, zero, or a positive real multiple of the imaginary unit. Assuming $w_{a}>0$, we have ${\pi_{a}{({x})}}=0$ iff $x$ lies on this sphere. Given two weighted points, ${\bf a}=(a,r_{a}^{2})$ and ${\bf b}=(b,r_{b}^{2})$ [^2] Because of the geometric picture with spheres, we find it more intuitive to talk about the radius rather than the weight, with the implicit understanding that the square of the radius can also be negative. with $a\neq b$, the bisector of $\bf a$} and $\bf b$} is the hyperplane of points $x\in{{\mathbb{R}}}^{d}$ that satisfy
${\mbox{$\pi_{a}{({x})}$}}-{\pi_{b}{({x})}}$ $=2{\langle b-a,x\rangle}-\left({\|{a}\|}^{2}-r_{a}^{2}-{\|{b}\|}^{2}+r_{b}^{2}\right)=0.$ (1)
It is normal to bab-a, but note that the point at which it intersects the line passing through $a$ and $b$ is not necessarily halfway between these points. While the bisector of two weighted points is a (d1)(d-1)-plane (a hyperplane), the set of points at equal power from $p+1\leq d+1$ weighted points is generically a (dp)(d-p)-plane. The only requirement for the existence of this (dp)(d-p)-plane is that the $p+1$ locations do not lie on a common (p1)(p-1)-plane. Specifically, for $p+1=d+1$ weighted points with affinely independent locations, this set is a single point.
[…]
Figure 4: Given two weighted points, ${\bf a}_{0}$ and ${\bf a}_{1}$, the circles of all weighted points orthogonal to both pass through the blue 0-sphere (pair of points) on the line that passes through $a_{0}$ and $a_{1}$.
We call ${\bf a}=(a,r_{a}^{2})$ and ${\bf b}=(b,r_{b}^{2})$ orthogonal to each other if ${\|{a}-{b}\|}^{2}=r_{a}^{2}+r_{b}^{2}$. By construction, for every point $x$ with equal power from $p+1\leq d+1$ weighted points, ${\bf a}_{0},{\bf a}_{1},\ldots,{\bf a}_{p}$, there is a unique weight, $w_{x}=r_{x}^{2}$, such that ${\bf{x}}=(x,r_{x}^{2})$ is orthogonal to each ${\bf a}_{i}$, for $0\leq i\leq p$. Let $z$ be the point that minimizes the common power from ${\bf a}_{0},{\bf a}_{1},\ldots,{\bf a}_{p}$, and note that it is the unique point at which the (dp)(d-p)-plane of points $x$ with equal power from $a_{0},a_{1},\ldots,a_{p}$ intersects the $p$-plane that passes through $a_{0},a_{1},\ldots,a_{p}$. Let $S$ be the (p1)(p-1)-sphere in which the sphere of ${\bf{z}}=(z,r_{z}^{2})$ intersects the $p$-plane. Importantly, the sphere of every weighted point orthogonal to each ${\bf a}_{i}$, for $0\leq i\leq p$, passes through $S$; see Figure 4 for an illustration for $p+1=2$ weighted points in ${{\mathbb{R}}}^{2}$. We formally state a generalization of this claim that allows for non-positive weights, and prove it for later reference.
Lemma 2.9.
Let ${\bf a}_{0},{\bf a}_{1},\ldots,{\bf a}_{p}$ be $p+1\leq d+1$ generic weighted points in ${{\mathbb{R}}}^{d}$, and ${\bf{z}}=(z,r_{z}^{2})$ the weighted point orthogonal to ${\bf a}_{i}$, for $0\leq i\leq p$, with minimum weight. Then ${\|{x}-{z}\|}^{2}=r_{x}^{2}-r_{z}^{2}$ for every weighted point ${\bf{x}}=(x,r_{x}^{2})$ that is orthogonal to ${\bf a}_{i}$ for $0\leq i\leq p$.
Proof 2.10.
By definition, $x$ and $z$ both lie on the (dp)(d-p)-plane of points with equal power from the $p+1$ weighted points. This (dp)(d-p)-plane is normal to bab-a, which implies that $x,z,a_{i}$ are the vertices of a right-angled triangle, so ${\|{x}-{a_{i}}\|}^{2}={\|{x}-{z}\|}^{2}+{\|{z}-{a_{i}}\|}^{2}$ for each $0\leq i\leq p$. To avoid double-indices, write $r_{i}^{2}$ for the weight of ${\bf a}_{i}$. Since $\bf{x}$} and $\bf{z}$} are both orthogonal to ${\bf a}_{i}$, we also have ${\|{x}-{a_{i}}\|}^{2}=r_{x}^{2}+r_{i}^{2}$ and ${\|{z}-{a_{i}}\|}^{2}=r_{z}^{2}+r_{i}^{2}$. Plugging the last two relations into the first, we get ${\|{x}-{z}\|}^{2}={\|{x}-{a_{0}}\|}^{2}-{\|{z}-{a_{0}}\|}^{2}=r_{x}^{2}-r_{z}^{2}$, as claimed.
We now return to main topic of this section. Writing $A\subseteq{{\mathbb{R}}}^{d}$ for the projection of ${\bf A}\subseteq{{\mathbb{R}}}^{d}\times{{\mathbb{R}}}$, we call $\bf A$} a thin Delone set of weighted points if $A$ is thin Delone and the absolute weights of the points are bounded away from $\infty$. Furthermore, $\bf A$} is generic if $A$ is generic and for every $d+1$ weighted points there is a unique and distinct weighted point that is orthogonal to each of the $d+1$ weighted points. The $d$-simplex defined by $d+1$ weighted points is the convex hull of their locations, and the heft of this simplex is the number of weighted points ${\bf b}=(b,r_{b}^{2})\in{\bf A}$ that satisfy ${\|{x}-{b}\|}^{2}-r_{x}^{2}-r_{b}^{2}$ ¡ 0, in which ${\bf{x}}=(x,r_{x}^{2})$ is the unique weighted point orthogonal to the $d+1$ weighted points.
@@ 2.5 Weighted Points
Let ${\bf A}\subseteq{{\mathbb{R}}}^{d}\times{{\mathbb{R}}}$ be a generic thin Delone set of weighted points, and $a$ the projection of a weighted point ${\bf a}\in{\bf A}$ to ${{\mathbb{R}}}^{d}$. Then the simplices with heft $k$ cover ${{\mathbb{R}}}^{d}$ exactly $\binom{d+k}{d}$ times, and the simplices with heft at most $k$ incident to $a$ cover every sufficiently small neighborhood of $a$ at most $\binom{d+k}{d}$ times.
Proof 2.12.
The proof of the global covering constant is similar to that of Theorem 2.3, and we present abridged versions of its three steps. Let $x\in{{\mathbb{R}}}^{d}$ be generic with respect to $A$, and write ${c}_{k}{({x,{\bf A}})}$} for the number of $k$-hefty simplices of $\bf A$} that contain $x$.
In Step I we show that ${c}_{k}{({x,{\bf A}})}$} does not depend on $x$. Indeed, if we move $x$ continuously it can possibly change only when $x$ crosses a facet of some $k$-hefty simplex, but even then ${c}_{k}{({x,{\bf A}})}$} is preserved, as we now show. Let $\Delta$ be such a facet and $L$ the line of points with equal power from the weighted points that project to the vertices of $\Delta$. For each generic point $y\in L$, mark $y$ with the number of weighted points ${\bf b}\in{\bf A}$ such that ${\pi_{b}{({y})}}<{\pi_{a}{({y})}}$ for each vertex $a$ of $\Delta$. Equivalently, ${\|{b}-{y}\|}<r_{b}^{2}+r_{y}^{2}$, in which ${\bf{y}}=(y,r_{y}^{2})$ is the unique sphere with center $y$ that is orthogonal to the spheres whose centers are the vertices of $\Delta$. This splits $L$ into open segments marked with integer values, and the number of $k$-hefty simplices that share $\Delta$ is the number of segments labeled $k$. Now use Lemma 2.9 while walking along $L$ from left to right. A $k$-hefty simplex with facet $\Delta$ lies to the left of $\Delta$ if it corresponds to a transition from a segment marked $k+1$ on the left to one marked $k$ on the right, and it lies to the right of $\Delta$ if the transition is the other way round. Since there are equally many transitions of either kind, there are equally many $k$-hefty simplices that share $\Delta$ on both sides. Hence, ${c}_{k}{({x,{\bf A}})}$} is independent of $x$.
In Step II we show that ${c}_{k}{({x,{\bf A}})}$} does not depend on $\bf A$} either, only on the dimension of the ambient space. This step requires a generalization of Lemma 2.1 to weighted points, which is easy because all absolute weights are bounded by a constant. We can therefore take bounded neighborhoods of two points, $x$ and $x^{\prime}$, place them without overlap next to each other, and prove ${{c}_{k}{({x,{\bf A}})}}={{c}_{k}{({x^{\prime},{\bf A}^{\prime}})}}$ as in the first step. In Step III we re-use the radial set of points displayed in Figure 2. The unweighted points are weighted points with weight 0, so we finally conclude that ${{c}_{k}{({x,{\bf A}})}}=\binom{d+k}{d}$, as claimed.
Finally, we prove the inequality for local covering. After removing $\bf a$} from $\bf A$}, the global covering constant proved above implies that for every $0\leq j\leq k-1$, exactly $\binom{d+j}{d}$ $j$-hefty simplices of ${\bf A}\setminus\{{\bf a}\}$ contain the point $a$. Adding $\bf a$} back, the heft of each of these simplices increases by at most $1$, so the number of simplices of heft at most $k$ that contain $a$ is at least
$M_{k-1}$ $=\tbinom{d+0}{d}+\tbinom{d+1}{d}+\ldots+\tbinom{d+k-1}{d}$. (2)
Indeed, it is also possible that a $k$-hefty simplex of ${\bf A}\setminus\{{\bf a}\}$ is only $k$-hefty for $\bf A$}, but this does not invalidate the derived inequality. The simplices of heft at most $k$ cover ${{\mathbb{R}}}^{d}$ exactly $M_{k}$ times, so the subset of simplices that are incident to $a$ cover any sufficiently small neighborhood of $a$ at most $M_{k}-M_{k-1}=\binom{d+k}{d}$ times.
The upper bound for local covering in Theorem 2.11 is tight, as we will see shortly. Indeed, if there were a smaller upper bound, then its application to proving an upper bound for $\Psi_{d}(k,n)$ in the next subsection would contradict the lower bound illustrated in Figure 6. In other words, that lower bound construction can be turned into a lower bound construction for the local covering question, and we leave the details to the interested reader.
3 Hefty Simplices on the Sphere
In this section, we extend the results of Section 2 from the Euclidean to the spherical setting. To a large extent, the proofs are similar to those in Section 2, so we limit ourselves to highlighting the differences.
3.1 (Unweighted) Points
The results in Theorems 2.3 and 2.5 can be extended to finite sets in ${{\mathbb{S}}}^{d}$. Note however that in the spherical setting, there is an ambiguity in the definition of a $k$-hefty simplex since any (d1)(d-1)-sphere on the $d$-sphere bounds two complementary $d$-balls. We thus require that the finite set satisfies the following condition.
Definition 3.1.
A set $A\subseteq{{\mathbb{S}}}^{d}$ is $k$-balanced if every open hemisphere contains at least $k+1$ points of $A$.
With this assumption, only the smaller of the two open $d$-balls has a chance to contain $k$ and no more of the points, so we consider it the inside of the (d1)(d-1)-sphere. The notions of genericity and heft are straightforward adaptions of those introduced in Euclidean space. In particular, a finite set $A\subseteq{{\mathbb{S}}}^{d}$ is generic if no $d+1$ points belong to a common (d1)(d-1)-dimensional great-sphere, and no $d+2$ points belong to a common (d1)(d-1)-dimensional sphere in ${{\mathbb{S}}}^{d}$. Equivalently, no $d+1$ of the points (unit vectors) are linearly dependent, and no $d+2$ are affinely dependent in ${{\mathbb{R}}}^{d+1}$. Furthermore, a point $a\in{{\mathbb{S}}}^{d}$ is generic with respect to $A$ if $A\cup\{a\}$ is generic. The heft of a spherical $d$-simplex spanned by $d+1$ points in $A$ is the number of points inside the unique (d1)(d-1)-sphere that passes through the $d+1$ points, and we call it a $k$-hefty simplex of $A$ if this number is $k$.
Theorem 3.2.